An ideal gas undergoes a mechanically reversible process (constant pressure, constant temperature and adiabatic process. The gas entering a T1=650K and P1=10bar decreases its temperature at constant pressure where V2=2.91x10- 3 m3. Then, it went to isothermal process to decrease its pressure. Finally, the gas returns to its initial state. Take

Cp=7

2


1
Expert's answer
2020-11-26T05:45:51-0500

T1=650K

P1= 10bar

V2=2.91x10-3

Constant P and T

V1=?

"Pdv= nRT"

10(V2-V1) = 8.314×650

2.91x10-2 -10v1 =5404.1

V1=5.4×10-3